# VideoPainter: Any-length Video Inpainting and Editing with Plug-and-Play Context Control

### v2

@@ 3.4 Plug-and-Play Control
4 Experiments
4.1 Implementation details
VideoPainter is built upon a pre-trained Image-to-Video Diffusion Transformer CogVideo-5B-I2V (Yang et al., 2024) (by default) and its Text-to-Video version. In training, we use VPData at a $480\times 720$ resolution, learning rate $1\times 10^{-5}$, batch size $1$ for both the context encoder ($80,000$ steps) and the ID Resample Adapter ($2,000$ steps) in two stages with AdamW on 64 NVIDIA V100 GPUs.
@@ 4.1 Implementation details
In video inpainting, we employ Davis (Perazzi et al., 2016) as the benchmark for random masks and VPBench for segmentation-based masks. VPBench consists of $100$ 6-second videos for standard video inpainting, and $16$ videos with an average duration of more than $30$ seconds for long video inpainting. The VPBench includes diverse content, including objects, humans, animals, landscapes, and multi-range masks. For video editing evaluation, we also utilize VPBench, which includes four fundamental editing operations (adding objects, removing objects, swaping objects, and changing attributes) and comprises $45$ 6-second videos and $9$ videos with an average duration of $30$ seconds.
@@
We consider $8$ key metrics from the following three aspects: masked region preservation, text alignment, and video generation quality.
@@
    Video Generation Quality. Following previous methods (Zhou et al., 2023), we use FVID (Wang et al., 2018) to measure the generated video quality.
4.2 Video Inpainting
Figure 5: Comparison of previous video inpainting methods and VideoPainter on standard and long video inpainting. More visualizations are in the demo video.
@@ 4.2 Video Inpainting
The qualitative comparison with previous video inpainting methods is shown in Fig. 5. VideoPainter consistently shows exceptional results in the video coherence, quality, and alignment with text caption. Notably, ProPainter fails to generate fully masked objects because it only depends on background pixel propagation instead of generating. While COCOCO demonstrates basic functionality, it fails to maintain consistent ID in inpainted regions ( inconsistent vessel appearances and abrupt terrain changes) due to its single-backbone architecture attempting to balance background preservation and foreground generation. Cog-Inp achieves basic inpainting results; however, its blending operation’s inability to detect mask boundaries leads to significant artifacts. Moreover, VideoPainter can generate coherent videos exceeding one minute while maintaining ID consistency through our ID resampling.
Table 2: Quantitative comparisons among VideoPainter and other video inpainting models in VPBench for segmentation mask (Standard (S) and Long (L) Video) and Davis for random mask: ProPainter (Zhou et al., 2023), COCOCO (Zi et al., 2024), and Cog-Inp (Yang et al., 2024). Metrics include masked region preservation, text alignment, and video quality. Red stands for the best, Blue stands for the second best.
@@
4.3 Video Editing
VideoPainter can be used for video inpainting by employing Vison Language Models (OpenAI, 2024; Team et al., 2024) to generate modified captions based on user editing instructions and source captions and apply VideoPainter to inpaint based on the modified captions. Tab. 3 shows the quantitative comparison on VPBench. We compare the editing results of inverse-based UniEdit (Bai et al., 2024), DiT-based DiTCtrl (Cai et al., 2024), and end-to-end ReVideo (Mou et al., 2024). For both standard and long videos in VPBench, VideoPainter achieves superior performance, even surpassing the end-to-end ReVideo. This success can be attributed to its dual-branch architecture, which ensures excellent background preservation and foreground generation capabilities, maintaining high fidelity in non-edited regions while ensuring edited regions closely align with editing instructions, complemented by inpainting region ID resampling that maintains ID consistency in long video. The qualitative comparison with previous video inpainting methods is shown in Fig. 5. VideoPainter demonstrates superior performance in preserving visual fidelity and text-prompt consistency. VideoPainter successfully generates a seamless animation of a futuristic spaceship traversing the sky, maintaining smooth temporal transitions and precise background boundaries throughout the removal process, without introducing artifacts that were observed in ReVideo.
Table 3: Quantitative comparisons among VideoPainter and other video editing models in VPBench (Standard and Long Video): UniEdit (Bai et al., 2024), DitCtrl (Cai et al., 2024), and ReVideo (Mou et al., 2024). Metrics include masked region preservation, text alignment, and video quality. Red stands for the best, Blue stands for the second best.
@@ 4.3 Video Editing
Figure 6: Comparison of previous video editing methods and VideoPainter on standard and long video editing. More visualizations are in the demo video.
Table 4: User Study: User preference ratios comparing VideoPainter with video inpainting and editing baselines.
@@ 4.4 Human Evaluation
4.5 Ablation Analysis
We ablate on VideoPainter in Tab .5, including framework architecture, context encoder size, group-wise feature integration strategy, and inpainting region ID resampling.
Table 5: Ablation Studies on VPBench. Single-Branch: We add input channels to adapt masked video and finetune the backbone. Layer Configuration (VideoPainter (*)): We vary the context encoder depth from one to four layers. w/o Selective Token Integration (w/o Select):: We bypass the token pre-selection step and integrate all context encoder tokens into DiT. T2V Backbone (VideoPainter (T2V)): We replace the backbone from image-to-video DiTs to text-to-video DiTs. w/o target region ID resampling (w/o Resample): We ablate on the target region ID resampling. (L) denotes evaluation on the long video subset. Red stands for the best result.
@@ 4.5 Ablation Analysis
Based on rows 1 and 5, the dual-branch VideoPainter significantly outperforms its single-branch counterpart by explicitly decoupling background preservation from foreground generation, thereby reducing model complexity and avoiding the trade-off between competing objectives in a single branch. Row 2 to row 6 of Tab. 5 demonstrate the rationale behind our key design choices: 1 utilizing a two-layer structure as an optimal balance between performance and efficiency for the context encoder, 2 implementing token-selective feature fusion based on segmentation mask information to prevent confusion from indistinguishable foreground-background tokens in the backbone, and 3 adapting plug-and-play control to different backbones with comparable performance. Furthermore, rows 7 and 8 verify the importance of employing inpainting region ID resampling for long videos, which preserves ID by explicitly resampling inpainted region tokens from previous clips.
4.6 Plug-and-Play Control Ability
@@
Figure 7: Integrating VideoPainter to Gromit-style LoRA (Cseti, 2024) in CogVideoX-5B (Yang et al., 2024).
5 Discussion
@@
In this paper, we introduce VideoPainter, the first dual-branch video inpainting framework with plug-and-play control capabilities. Our approach features three key innovations: (1) a lightweight plug-and-play context encoder compatible with any pre-trained video DiTs, (2) an inpainting region ID resampling technique for maintaining long video ID consistency, and (3) a scalable dataset pipeline that produced VPData and VPBench, containing over 390K video clips with precise masks and dense captions. VideoPainter also shows promise in video editing applications. Extensive experiments demonstrate that VideoPainter achieves state-of-the-art performance across 8 key metrics in video inpainting and video editing, particularly in video quality, mask region preservation, and text coherence.
@@
However, VideoPainter still has several limitations: (1) Generation quality is limited by the base video DiT models, which may struggle with complex physical and motion modeling scenarios, and (2) performance is suboptimal with low-quality masks or misaligned video captions.

### v3

@@ 3.4. Plug-and-Play Control
4. Experiments
4.1. Implementation details
VideoPainter is built upon a pre-trained Image-to-Video Diffusion Transformer CogVideo-5B-I2V (Yang et al., 2024) (by default) and its Text-to-Video version. In training, we use VPData at a $480\times 720$ resolution, learning rate $1\times 10^{-5}$, batch size $1$ for both the context encoder ($80,000$ steps) and the ID Resample Adapter ($2,000$ steps) in two stages with AdamW. In training, we randomly sample dilation and erosion with kernel sizes $\in[8,32]$ to enhance robustness to mask precision. This also enables our random-mask inpainting.
@@ 4.1. Implementation details
In video inpainting, we employ Davis (Perazzi et al., 2016) as the benchmark for random masks and VPBench for segmentation-based masks. VPBench consists of $100$ 6-second videos for standard video inpainting, and $16$ videos with an average duration of more than $30$ seconds for long video inpainting. The VPBench includes diverse content including objects, humans, animals, landscapes, and multi-range masks. For video editing evaluation, we also utilize VPBench, which includes four fundamental editing operations (add, remove, swap, and change) and comprises $45$ 6-second videos and $9$ videos with an average duration of $30$ seconds.
@@
We consider $8$ metrics from three aspects: masked region preservation, text alignment, and video generation quality.
@@
    Video Generation Quality. Following previous methods, we use FVID (Wang et al., 2018) to measure the generated video quality.
4.2. Video Inpainting
Figure 5. Comparison of previous inpainting methods and VideoPainter on standard and long video inpainting. More visualizations are in the demo video.
@@ 4.2. Video Inpainting
The qualitative comparison with previous video inpainting methods is shown in Fig. 5. VideoPainter consistently shows exceptional results in video coherence, quality, and alignment with text caption. Notably, ProPainter fails to generate fully masked objects because it only depends on background pixel propagation instead of generating. While COCOCO demonstrates basic functionality, it fails to maintain consistent ID in inpainted regions ( inconsistent vessel appearances and abrupt terrain changes) due to its single-backbone architecture attempting to balance background preservation and foreground generation. Cog-Inp achieves basic inpainting results; however, its blending operation’s inability to detect mask boundaries leads to significant artifacts. Moreover, VideoPainter can generate coherent videos exceeding one minute while maintaining ID consistency through our ID resampling.
Table 2. Quantitative comparisons among VideoPainter and other video inpainting models in VPBench for segmentation mask (Standard (S) and Long (L) Video) and Davis for random mask: ProPainter (Zhou et al., 2023), COCOCO (Zi et al., 2024), and Cog-Inp (Yang et al., 2024). Metrics include masked region preservation, text alignment, and video quality. Red stands for the best, Blue stands for the second best.
@@
4.3. Video Editing
VideoPainter can be used for video inpainting by employing Vison Language Models (OpenAI, 2024; Team et al., 2024) to generate modified captions based on user editing instructions and source captions and apply VideoPainter to inpaint based on the modified captions. Tab. 3 shows the quantitative comparison on VPBench. We compare the editing results of inverse-based UniEdit (Bai et al., 2024), DiT-based DiTCtrl (Cai et al., 2024), and end-to-end ReVideo (Mou et al., 2024). For both standard and long videos in VPBench, VideoPainter achieves superior performance, even surpassing the end-to-end ReVideo. This success can be attributed to its dual-branch architecture, which ensures excellent background preservation and foreground generation capabilities, maintaining high fidelity in non-edited regions while ensuring edited regions closely align with editing instructions, complemented by inpainting region ID resampling that maintains ID consistency in long video. The qualitative comparison with previous video inpainting methods is shown in Fig. 5. VideoPainter demonstrates superior performance in preserving visual fidelity and text-prompt consistency.
Table 3. Quantitative comparisons among VideoPainter and other video editing models in VPBench (Standard and Long Video): UniEdit (Bai et al., 2024), DitCtrl (Cai et al., 2024), and ReVideo (Mou et al., 2024). Metrics include masked region preservation, text alignment, and video quality. Red stands for the best, Blue stands for the second best.
@@ 4.3. Video Editing
Figure 6. Comparison of previous editing methods and VideoPainter on standard and long video editing. More visualizations are in the demo video.
Table 4. User Study: User preference ratios comparing VideoPainter with video inpainting and editing baselines. For each sample, participants selected only one model that produced the best results for each criterion. We evaluate performance using the average proportion of being selected as the best response. For video inpainting, we compared VideoPainter against ProPainter (Zhou et al., 2023), COCOCO (Zi et al., 2024), and Cog-Inp (Yang et al., 2024). For video editing, we compared VideoPainter against UniEdit (Bai et al., 2024), DitCtrl (Cai et al., 2024), and ReVideo (Mou et al., 2024). Detailed results are in the appendix.
@@ 4.4. Human Evaluation
4.5. Ablation Analysis
We ablate on VideoPainter in Tab .7, including architecture, context encoder size, control strategy, and inpainting region ID resampling.
Table 5. Ablation Studies on VPBench. Single-Branch: We add input channels to adapt masked video and finetune the backbone. Layer Configuration (VideoPainter (*)): We vary the context encoder depth from one to four layers. w/o Selective Token Integration (w/o Select):: We bypass the token pre-selection step and integrate all context encoder tokens into DiT. T2V Backbone (VideoPainter (T2V)): We replace the backbone from image-to-video DiTs to text-to-video DiTs. w/o target region ID resampling (w/o Resample): We ablate on the target region ID resampling. (L) denotes evaluation on the long video subset. Red stands for the best result.
@@ 4.5. Ablation Analysis
Based on rows 1 and 5, the dual-branch VideoPainter significantly outperforms its single-branch counterpart by explicitly decoupling background preservation from foreground generation, thereby reducing model complexity and avoiding the trade-off between competing objectives in a single branch. Row 2 to row 6 of Tab. 7 demonstrate the rationale behind our key design choices: 1 utilizing a two-layer structure as an optimal balance between performance and efficiency for the context encoder, 2 implementing token-selective feature fusion based on segmentation mask information to prevent confusion from indistinguishable foreground-background tokens in the backbone, and 3 adapting plug-and-play control to different backbones with comparable performance. Furthermore, rows 7 and 8 verify the importance of employing inpainting region ID resampling for long videos, which maintains ID consistency by explicitly resampling inpainted region tokens from previous clips.
4.6. Plug-and-Play Control Ability
@@
Figure 7. Integrating VideoPainter to Gromit-style LoRA (Cseti, 2024).
5. Discussion
@@
In this paper, we introduce VideoPainter, the first dual-branch video inpainting framework with plug-and-play control capabilities. Our approach features three key innovations: (1) a lightweight plug-and-play context encoder compatible with any pre-trained video DiTs, (2) an inpainting region ID resampling technique for maintaining long video ID consistency, and (3) a scalable dataset pipeline that produced VPData and VPBench, containing over 390K video clips with precise masks and dense captions. VideoPainter also shows promise in video editing applications. Extensive experiments demonstrate that VideoPainter achieves state-of-the-art performance across 8 metrics in video inpainting and editing, particularly in video quality, mask region preservation, and text coherence.
@@
However, VideoPainter still has limitations: (1) Generation quality is limited by the base model, which may struggle with complex physical and motion modeling, and (2) performance is suboptimal with low-quality masks or misaligned video captions.
Acknowledgements
This work was supported in part by the CUHK Strategic Seed Funding for Collaborative Research Scheme under Grant No. 3136023 and in part by the CUHK Research Matching Scheme under Grant No. 7106937, 8601130, and 8601440.
Appendix A User Study
For a comprehensive evaluation beyond the quantitative and qualitative experiments presented in sections 4.2 and 4.3, we conducted a user study on both video inpainting and editing tasks using standard-length video samples from the VPBench inpainting subset and editing subset. The study involved 30 participants evaluating 50 randomly selected cases across three criteria: background preservation, text alignment, and video quality. Specifically, we assessed (1) Background Preservation: the error distance between the uninpainted regions of the generated video and the original video, (2) Text Alignment: the semantic coherence between inpainted regions and text captions, and (3) Video Quality: the overall visual fidelity of the generated results. As shown in Tab. 6, our approach demonstrates superior performance across all three aspects in both tasks. For video inpainting, our method achieves win rates of 74.2%, 82.5%, and 87.4% respectively. Similarly, for video editing, we obtain win rates of 78.4%, 76.1%, and 81.7% across the three aspects.
Table 6. User study evaluation comparing VideoPainter against state-of-the-art video inpainting and editing models. We conducted comprehensive comparisons on the VPBench, randomly sampling 50 examples from each of the inpainting and editing subsets. Human evaluators assessed the models’ outputs based on three criteria: background preservation, text alignment, and overall video quality. For each sample, participants selected only one model that produced the best results for each criterion. We evaluate performance using the proportion of model-generated outputs selected as the optimal response across all samples. For video inpainting, we compared against ProPainter (Zhou et al., 2023), COCOCO (Zi et al., 2024), and Cog-Inp (Yang et al., 2024). For video editing, we compared against UniEdit (Bai et al., 2024), DitCtrl (Cai et al., 2024), and ReVideo (Mou et al., 2024). Red stands for the best, Blue stands for the second best.
Metrics Background Text Video
[…]
Inpainting ProPainter 3.5 2.5 1.7
COCOCO 7.0 2.7 2.1
Cog-Inp 15.3 12.3 8.8
Ours 74.2 82.5 87.4
Editing UniEdit 3.3 3.7 2.8
DitCtrl 5.8 6.1 4.2
ReVideo 12.5 14.1 11.3
Ours 78.4 76.1 81.7
Appendix B Ablation on Dual-Branch Artecture
Table 7. Ablation Studies on VPBench. Single-Branch: We add input channels to adapt masked video and finetune the backbone. Layer Configuration (VideoPainter (*)): We vary the context encoder depth from one to four layers. w/o Selective Token Integration (w/o Select):: We bypass the token pre-selection step and integrate all context encoder tokens into DiT. T2V Backbone (VideoPainter (T2V)): We replace the backbone from image-to-video DiTs to text-to-video DiTs. w/o target region ID resampling (w/o Resample): We ablate on the target region ID resampling. (L) denotes evaluation on the long video subset. Red stands for the best result.
Metrics Masked Region Preservation Text Alignment Video Quality
[…]
Single-Branch 20.54 0.79 10.48 0.94 4.16 8.19 19.31 0.22
VideoPainter (1) 21.92 0.81 8.78 0.89 3.26 8.44 20.79 0.17
VideoPainter (4) 22.86 0.85 6.51 0.83 2.86 9.12 20.49 0.16
w/o Select 20.94 0.74 7.90 0.95 3.87 8.26 17.84 0.25
VideoPainter (T2V) 23.01 0.87 6.94 0.89 2.65 9.41 20.66 0.16
VideoPainter 23.32 0.89 6.85 0.82 2.62 8.66 21.49 0.15
w/o Resample (L) 21.79 0.84 8.65 0.81 3.10 11.35 20.68 0.19
VideoPainter (L) 22.19 0.85 9.14 0.71 2.92 11.52 21.54 0.17
Figure 8. Training loss curve of ablation of single branch fine-tuning and default dual branch VideoPainter. The training loss curves demonstrate that our dual-branch VideoPainter achieves superior convergence speed, stability, and final performance compared to single-branch fine-tuning, despite having significantly fewer trainable parameters.
As shown in Tab. LABEL:tab:ablation_on_branch and Fig. 8, we ablate on VideoPainter with a single branch fine-tuning setup.
Tab. LABEL:tab:ablation_on_branch clearly shows that the dual-branch VideoPainter significantly outperforms its single-branch counterpart by explicitly decoupling background preservation from foreground generation, thereby reducing model complexity and avoiding the trade-off between competing objectives in a single branch. The training loss curves in Fig. 8 demonstrate that VideoPainter, through effective dual-branch decoupling, achieves an advantageous optimization starting point and maintains stable, efficient convergence. By separating the competing objectives of background preservation and foreground generation via follow caption into distinct branches rather than a single backbone, VideoPainter achieves superior convergence performance, with a final loss approximately half that of single-branch finetuning.
Appendix C Ablation on Mask Quality
Table 8. Ablation Studies on VPBench: Kernel (*): We randomly sample dilation and erosion with varying kernel sizes $\in(8,16,32)$ for the segmentation masks. Kernel (Square): We randomly sample square masks with varying sizes $\in[8,32]$ and random locations. This reflects VideoPainter’s robustness to different mask qualities. Red stands for the best result.
Metrics Masked Region Preservation Text Alignment Video Quality
[…]
Kernel (8) 23.24 0.84 7.43 0.84 2.69 8.51 21.18 0.15
Kernel (16) 23.17 0.82 7.55 0.87 2.93 8.49 20.37 0.17
Kernel (32) 23.02 0.81 7.64 0.92 3.07 8.28 20.38 0.16
Kernel (Square) 22.46 0.73 9.37 0.94 3.75 7.81 19.88 0.21
VideoPainter 23.32 0.89 6.85 0.82 2.62 8.66 21.49 0.15
As shown in Tab. 8, we perform ablation studies on VideoPainter using a single branch fine-tuning configuration.
Tab. 8 clearly demonstrates that VideoPainter exhibits robust generalization across various qualities of video segmentation masks (which we augmented by randomly sampling different kernel sizes for dilation and erosion operations applied to the original segmentation masks). However, when utilizing random rectangular masks (which are highly atypical in video segmentation), model performance decreased by approximately 13%. This performance degradation primarily stems from artifacts generated along the rectangular edges, which we attribute to the substantial geometric disparity between these artificial rectangular masks and natural segmentation masks encountered in real-world scenarios.
Appendix D Ablation on ID Resampling
Table 9. Ablation Studies on VPBench: w/o target region ID resampling (w/o Resample): We ablate on the target region ID resampling. (L) denotes evaluation on the long video subset. Red stands for the best result.
Metrics Masked Region Preservation Text Alignment Video Quality
[…]
w/o Resample (L) 21.79 0.84 8.65 0.81 3.10 11.35 20.68 0.19
VideoPainter (L) 22.19 0.85 9.14 0.71 2.92 11.52 21.54 0.17
Figure 9. Long video inpainting comparison of original VideoPainter and w/o ID resampling ablation version.
As shown in Tab. 9 and Fig. 9, we ablate on VideoPainter with a w/o ID resampling ablation version.
Tab. LABEL:tab:ablation_on_branch clearly verifies the importance of employing inpainting region ID resampling for long videos, which maintains ID consistency by explicitly resampling inpainted region tokens from previous clips. Fig. 8 demonstrates that removing the ID resample adapter degrades long-video inpainting performance, with more severe deterioration as video length increases.
Appendix E More Related Works
E.1. Video Inpainting
Video inpainting approaches can be broadly classified into two categories based on whether they possess generative capabilities:
Non-generative methods.
These methods (Zhou et al., 2023; Li et al., 2022; Hu et al., 2020; Zhang et al., 2022a, b) leverage architecture priors to facilitate pixel propagation. This includes utilizing local perception of 3D CNNs (Chang et al., 2019a; Wang et al., 2019; Hu et al., 2020; Chang et al., 2019b), and exploiting the global perception of attention to retrieve and aggregate tokens with similar texture for filling masked video (Lee et al., 2019; Zeng et al., 2020; Liu et al., 2021; Zhang et al., 2022a). They also introduce various physical quantities, especially optical flow, as auxiliary conditions as it simplifies RGB pixel inpainting by completing less complex flow fields (Kim et al., 2019; Li et al., 2020; Zou et al., 2021; Xu et al., 2019; Gao et al., 2020; Zhang et al., 2022c, b). A notable exemplar of this approach is ProPainter (Zhou et al., 2023), which integrates flow completion models with spatiotemporal transformers into an end-to-end framework, achieving more faithful propagation through combined pixel and feature propagation with flow consistency verification. However, they are only effective for partial object occlusions with random masks but face significant limitations when inpaint fully masked regions due to insufficient contexts.
Generative methods.
Recent advances in generative foundation models (Rombach et al., 2022; Guo et al., 2023) have sparked numerous approaches that leverage additional modules or training strategies to extend backbones’ capabilities for video inpainting (Zhang et al., 2024b; Zi et al., 2024; Wang et al., 2024). VideoComposer (Wang et al., 2024), built upon a pre-trained video generation backbone Stable Video Diffusion (Blattmann et al., 2023), integrates various control signals (text, depth, mask, motion vectors) through a shared spatiotemporal condition fusion module. While it enables limited inpainting capabilities through mask and text conditioning, its performance in background preservation is compromised due to control condition compression and lack of masked video conditioning support. AVID (Zhang et al., 2024b) and COCOCO (Zi et al., 2024) represent the most related recent works. Both adopt a similar implementation by augmenting Stable Diffusion Inpainting (Rombach and Esser, 2022) with trainable temporal attention layers. This architecture includes per-frame region filling based on the image inpainting backbone and temporal smoothing with temporal attention. Despite showing promising results for both random and segmentation masks due to their generative abilities, they struggle to balance background preservation and foreground generation with text caption (Ju et al., 2024; Li et al., 2024; Bian et al., 2024b) within the single backbone. AVID also explores any-length video inpainting by smoothing latent at segment boundaries and using the middle frame as the ID reference. However, identity discontinuities occur during significant variations, and identity degradation emerges in longer video generation scenarios. In contrast, VideoPainter is a dual-branch framework by decoupling video inpainting into foreground generation and background-guided preservation. It employs an efficient context encoder to guide any pre-trained DiT, facilitating plug-and-play control and zero-shot adaptation across various stylization backbones—a capability absent in both non-generative and generative methods. Furthermore, VideoPainter also introduces a novel target region ID resampling technique that enables ID consistency in any-length video inpainting by explicitly resampling the inpainted region of the last generated clip.
Table 10. Comparison of video inpainting datasets. Our VPData is the largest video inpainting dataset to date, comprising over 390K high-quality clips with segmentation masks, video captions, and masked region descriptions.
Dataset #Clips Duration Video Caption Masked Region Desc.
DAVIS (Perazzi et al., 2016) 0.4K 0.1h
YouTube-VOS (Xu et al., 2018) 4.5K 5.6h
VOST (Tokmakov et al., 2023) 1.5K 4.2h
MOSE (Ding et al., 2023) 5.2K 7.4h
LVOS (Hong et al., 2023) 1.0K 18.9h
SA-V (Ravi et al., 2024) 642.6K 196.0h
Ours 390.3K 866.7h
E.2. Video Inpainting Datasets
Recent advances in segmentation (Ravi et al., 2024) have created many video segmentation datasets (Perazzi et al., 2016; Xu et al., 2018; Hong et al., 2023; Ding et al., 2023; Tokmakov et al., 2023; Darkhalil et al., 2022). Among these, DAVIS (Perazzi et al., 2016) and YouTube-VOS (Xu et al., 2018) have become prominent benchmarks for video inpainting due to their high-quality masks and diverse object categories. However, the existing datasets face two primary limitations: (1) insufficient scale to meet the data requirements of generative models, and (2) the absence of crucial control conditions necessary for generating masked objects such as video captions. These constraints have impeded the evaluation of state-of-the-art generative approaches, with methods like COCOCO and AVID resorting to random sampling from annotated internet video datasets for assessment. Consequently, these limitations have also hindered the further advancement of generative approaches in video inpainting. In contrast, as shown in Tab. 10, we developed a scalable dataset pipeline based on state-of-the-art vision understanding models (Zhang et al., 2024a; Ravi et al., 2024; OpenAI, 2024), and constructed the largest video inpainting dataset to date with over 390K clips, each annotated with segmentation masks and dense video captions. In contrast, we developed a scalable dataset construction pipeline based on state-of-the-art vision understanding models (Zhang et al., 2024a; Ravi et al., 2024; OpenAI, 2024), including open-set object detection, object segmentation mask tracking, shot transition detection, motion score and aesthetic quality filtering, and video annotation. By leveraging publicly available internet videos, we constructed the largest video inpainting dataset to date, comprising over 390K high-quality clips, each accompanied by precise segmentation masks, comprehensive video captions, and masked region descriptions, as Tab. 10 shows.
Appendix F More Discussions
As mentioned in the main paper, VideoPainter’s performance is suboptimal with low-quality or misaligned video captions. However, it should be noted that VideoPainter’s text-following capability (including complex physical or motion modeling abilities) derives from pre-trained DiTs. This can be enhanced by substituting more powerful DiT models, as our method is a plug-and-play framework independent of the internal DiT backbone architecture.
Appendix G More Visualizations
Figure 10. More video inpainting results.
Figure 11. More video editing results.